OpenStudy (anonymous):

simplify by factoring. Assume that all expressions under radicals represent nonneggative numbers

OpenStudy (anonymous):

\[\sqrt[3]{189x^8}\]

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{3\times 3 \times 7 x^8}\]

OpenStudy (anonymous):

\[\sqrt[3]{63x^8}\]

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

Ops, i missed a 3.

OpenStudy (saifoo.khan):

\[\Large \sqrt[3]{3 \times 3 \times 3 \times 7x^8}\]

OpenStudy (anonymous):

189?

OpenStudy (saifoo.khan):

\[\Large 3 x^2 \sqrt[3]{7 x^2}\]

OpenStudy (anonymous):

ok I was way off :)

OpenStudy (saifoo.khan):

;)